Two forces of magnitude 80𝑁 and 45𝑁 act on an object, with an angle between the forces of 100Β°. What is the resultant force?


1
Expert's answer
2021-05-03T05:05:08-0400

Let force "F_1=80N"

Let the force "F_2=45N"

The resultant of the two forces can be obtained using:


"R^2 = F_1^2+F_2^2-2F_1F_2\cos\theta\\"


The angle formed between the two forces lies in the second quadrant. Thus the value of "\cos\theta" will be negative.

Thus:


"R^2 = 80^2 + 45^2 - 2\times80\times45\times\cos(180\degree-100\degree)\\
R^2 = 6400 + 2025 -7200 \times \cos80\degree\\
R^2 = 8425 - 7200 \times 0.1737\\
R^2 = 8425 -1,250.64\\
R^2 = 7,174.36\\
R = \sqrt{7,174.36}\\
R = 84.7N"